What is the heat on GDP a $100 billion charge in planned investment the MPC is 0.75? _______ billion.
Answer #1

Multiplier =1/(1-MPC)
=1/(1-0.75)
=4

change in real GDP=change in investment * multiplier
=100*4
=$400 billion